OpenStudy (anonymous):

use the discriminiant to determine the nature of the roots of x^2-4x+4=0

OpenStudy (gorv):

discriminiant=(-4)^2-4*1*4=26-16=0

OpenStudy (gorv):

sorry 16-16=0

OpenStudy (gorv):

so root will be real n equal

OpenStudy (gorv):

@grisell95